OpenStudy (anonymous):

A=40 degrees, C=75 degrees, c=20 does it make one triangle, two triangles, or no triangle?

OpenStudy (anonymous):

Do you know total degree in a triangle?

OpenStudy (anonymous):

yes its 180 which mekes angle B=65 degrees

OpenStudy (anonymous):

then I use law of Sines to get b/Sin65=20/sin75

OpenStudy (anonymous):

please give me a sec to put this in my calc

OpenStudy (anonymous):

alright So I got b=18.55

OpenStudy (anonymous):

So do i just do like the same thing to get A?

OpenStudy (dumbcow):

yes and from there you could solve for a ....but this only gives you 1 triangle with 3 distinct angles/sides

OpenStudy (anonymous):

So how exactly do you know if it equals one triangle or two?

OpenStudy (anonymous):

WaIT! so do I do C/sin c= A/sin a or B/sin b= A/ sin a?

OpenStudy (dumbcow):

i believe its from solving for the angles...sometimes you may get a situation where when solving for an angle using Law of sines, there will be 2 possible answers example \[\frac{\sin A}{3} = \frac{\sin 30}{4}\] \[A = \sin^{-1} (\frac{3\sin 30}{4}) = \sin^{-1} (\frac{3}{8})\]

ganeshie8 (ganeshie8):

also, by AAS congruence theorem, only one triangle is possible given a set of fixed Angle Angle and Side (40, 75, 20)

OpenStudy (dumbcow):

yes to find any side or angle, you can use Law of Sines \[\frac{a}{\sin A} = \frac{b}{\sin B}= \frac{c}{\sin C}\] or \[\frac{\sin A}{a} = \frac{\sin B}{b} = \frac{\sin C}{c}\] doesn't matter

OpenStudy (anonymous):

okay so now I just want to know how you know for sure if its one triangle or two or none

OpenStudy (dumbcow):

look at example i did in earlier post...notice there will be 2 solutions for angle A this means there are two possible triangles

ganeshie8 (ganeshie8):

alternative of law of sines you can use, congruency theorems + triangle inequality theorem : congruency theorems : AAS : exactly one triangle ASA : exactly one triangle SAS : exactly one triangle SSS : exactly one triangle apply triangle inequality theorem for below : SSA : exactly two triangles rest : infinity/none

OpenStudy (anonymous):

Alright so I have one saying it's one triangle and another saying it's two, so Who's right?

ganeshie8 (ganeshie8):

dumbcow is referring to previous post.... not this one

ganeshie8 (ganeshie8):

angle A is given here,

OpenStudy (anonymous):

oh okay so how did he get two answers? Cuz I don't get how he got sin^-1(3/8)

ganeshie8 (ganeshie8):

hmm im not following law of sines.... im for congruency theorems, they are more intuitive and look simple to me

OpenStudy (anonymous):

okay yeah that looks simpler to me too except for how SAS is one triangle while SSA is two

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

if SIDE, ANGLE and SIDE are fixed in that order, its obvious that only one triangle cuz, you cant do anything here, except connecting the two segment ends, right ?

OpenStudy (anonymous):

yeah So if I has SSA how do you get two?

ganeshie8 (ganeshie8):